OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find f(4).

OpenStudy (yrelhan4):

Just substitute 4 for x in f(x), which means.. f(4)=2*4 -4..

OpenStudy (anonymous):

Would that be f=1?

OpenStudy (yrelhan4):

solve this.. (2*4)-4, where * sign means multiplication..

OpenStudy (anonymous):

f=4?

OpenStudy (yrelhan4):

yup. f(4)=4..

OpenStudy (imtiaz7):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find f(4). i'll solve it for you @Nikkib312 for f(x) if f=4 then put it: f(x) = 2x-4 f(4) = 2(4)-4 f(4) =8-4 f(4) =4 so what about g(x)?

OpenStudy (yrelhan4):

You are not supposed to give direct answers. Its against the COC.

OpenStudy (imtiaz7):

what is coc

OpenStudy (yrelhan4):

Code of Conduct. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Thank you